OpenStudy (anonymous):

HELP :) MEDALS _THe perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm2.What are the dimensions of the rectangle ? a.11 cm by 16 cm b.8cm by 22 cm c.5.5 cm by 32 cm d.4cm by 44 cm

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@wwhitlock can you show me how to solve this ?

OpenStudy (anonymous):

p = 2*l + 2*w p = 54 a = l*w a = 176 2 equations, can you see what the next step is?

OpenStudy (anonymous):

Hi no I need help

OpenStudy (anonymous):